OpenStudy (anonymous):

What is the derivative of e^(x+3lnx)

OpenStudy (anonymous):

The derivative of e^u is e^u du The derivative of e^(x + 3 ln x) is e^(x + 3 ln x) times (1 + 3/x) Final

OpenStudy (jhannybean):

You got it :)

OpenStudy (jhannybean):

Let's say \(\large x+3\ln(x) = f(x)\), then we would have \(\large e^{f(x)}\) \[\large \frac{d}{dx} e^{f(x)} =f'(x) e^{f(x)} \]